OpenStudy (anonymous):

1. –6p + 7 = 3(2p – 3) – 4(–10 + 4p) p = 6 p = 5 p = 7 p = 12 I can't figure this out,i tried all but they don't equal the same on both sides.

OpenStudy (aroub):

p = 6 because, -6p+7=6p-9+40-16p -6p-6p+16p=-7-9+40 4p=24 p=6 :D

OpenStudy (ashish56):

p=6